OpenStudy (anonymous):

Find solutions for sinθ=cos2θ

OpenStudy (anonymous):

A solution using Mathematica 9 is attached.

OpenStudy (anonymous):

Given sinθ=cos2θ , we have Left= \[\sin \theta=\cos ^{2} \theta-\sin ^{2} \theta=1-2\sin ^{2} \theta\] =Right, So, let x=sinθ, \[2x ^{2}+x-1=0\] x1=-1,x2=1/2; So, 1. If sinθ=-1, \[\theta=\frac{ 3 }{ 2 }θ+2(\pi)k, k=Z;\] 2. If sinθ=1/2, \[\theta=\frac{1 }{4 }θ+2(\pi)k, k=Z;\]